# Exploiting Vocabulary Frequency Imbalance in Language Model Pre-training

### v2

@@ Appendix B Coverage of the most frequent N words
In this section, we measure the coverage of the frequent words in Fineweb-edu [40] and OpenWebText [13]. Both dataset exhibit a steep rise in cumulative coverage as we include more high-frequency tokens, but with subtly different baselines and slopes. In figure 6(a), the most frequent $500$ words already cover about 5858% of all tokens, climbing to roughly 7575% once we take the $2,500$ most frequent words. OpenWebText (Figure 6(b)) starts marginally higher—around 5959% at most frequent $500$ words, but follows an almost identical trajectory, reaching about 7676% coverage by the frequent $2,500$ words. This pattern underscores how a relatively small core vocabulary captures the vast majority of running text in both corpora, with only modest gains as we move deeper into the long tail.
@@ Appendix D OLMo-2 result
To identify whether the reduction in frequent words loss with increasing model size holds for contemporary large language models, we perform analogous experiments using the OLMo-2 series [38]. Figure 8(a) and 8(b) indicate that the average per vocabulary loss falls from $3.39$ nats in the 11B parameter model to $3.07$ nats in the 77B variant while slightly increasing the proportion of loss covered by $2,500$ frequent words from $71\%$ to $72.5\%$. Figure 8(b) further shows that global cross-entropy loss declines from $2.35$ nats for OLMo-2 $1\mathrm{B}$ to $2.12$ nats for OLMo-2 77B. Notably, OLMo-2 employs a much larger vocabulary (cl100K [39]) than Pythia (50304 tokens [6]) and trains on a larger corpus [38], which helps drive down the average loss on high-frequency words. These results confirm that the same trend holds for modern large-scale language models as well.
@@ Appendix E Are the gains mainly driven by parameter increases from vocab expansion?
To probe whether gains come from parameter growth with larger vocabularies, we run a controlled study comparing a $24\mathrm{K}$ vocabulary model ($122\mathrm{M}$ parameters) to a near equal size $49\mathrm{K}$ model ($124\mathrm{M}$ parameters) to isolate embedding capacity effects. We reduced the hidden dimension from $768$ to $648$ and the feed‑forward dimension from $2048$ to $1728$, while maintaining identical model depth which known to have a dominant impact on performance [56, 41]. As shown in the table 3, the $49\mathrm{K}$ model incurs higher cross‑entropy loss and delivers inferior results on downstream benchmarks (ARC‑E, HellaSWAG, PIQA, and SCIQ) compared to its $24\mathrm{K}$ counterpart, indicating that simply expanding the embedding layer contributes far less to model expressivity than is often assumed.
@@ Appendix F Do the benefits of larger vocabularies persist with tied-embedding models?
Several studies show that tying input and output embeddings can change pre-training dynamics [22, 29]. We therefore repeat our study with tied embeddings and report average frequent-word loss and cross-entropy on FineWeb-Edu in the table 4. The tied embedding variant shows broadly similar average frequent word loss and ce-loss to the untied baseline.
@@ Appendix H Learning rate exploration
We use a fixed learning rate of $6\times 10^{-4}$ for an $85\mathrm{M}$ non-embedding–parameter language model, mirroring the learning rate of GPT-3 with an equivalent non-embedding model size [7]. we further conducted a learning‑rate exploration with four additional learning rates: $2.4\times 10^{-3}$, $1.2\times 10^{-3}$, $1.5\times 10^{-4}$, $7.5\times 10^{-5}$. Table 6 and 7 report the resulting cross‑entropy loss and average word loss of frequent 2500 words in fineweb-edu across different vocabulary sizes and learning rates, using the same validation dataset described in Section 3.1.

### v3

@@ Appendix B Coverage of the most frequent N words
In this section, we measure the coverage of the frequent words in Fineweb-edu [penedo2024finewebdatasetsdecantingweb] and OpenWebText [Gokaslan2019OpenWeb]. Both dataset exhibit a steep rise in cumulative coverage as we include more high-frequency tokens, but with subtly different baselines and slopes. In figure 6(a), the most frequent $500$ words already cover about 5858% of all tokens, climbing to roughly 7575% once we take the $2,500$ most frequent words. OpenWebText (Figure 6(b)) starts marginally higher—around 5959% at most frequent $500$ words, but follows an almost identical trajectory, reaching about 7676% coverage by the frequent $2,500$ words. This pattern underscores how a relatively small core vocabulary captures the vast majority of running text in both corpora, with only modest gains as we move deeper into the long tail.
@@ Appendix D OLMo-2 result
To identify whether the reduction in frequent words loss with increasing model size holds for contemporary large language models, we perform analogous experiments using the OLMo-2 series [olmo2]. Figure 8(a) and 8(b) indicate that the average per vocabulary loss falls from $3.39$ nats in the 11B parameter model to $3.07$ nats in the 77B variant while slightly increasing the proportion of loss covered by $2,500$ frequent words from $71\%$ to $72.5\%$. Figure 8(b) further shows that global cross-entropy loss declines from $2.35$ nats for OLMo-2 $1\mathrm{B}$ to $2.12$ nats for OLMo-2 77B. Notably, OLMo-2 employs a much larger vocabulary (cl100K [gpt-4]) than Pythia (50304 tokens [gpt-neox]) and trains on a larger corpus [olmo2], which helps drive down the average loss on high-frequency words. These results confirm that the same trend holds for modern large-scale language models as well.
@@ Appendix E Are the gains mainly driven by parameter increases from vocab expansion?
To probe whether gains come from parameter growth with larger vocabularies, we run a controlled study comparing a $24\mathrm{K}$ vocabulary model ($122\mathrm{M}$ parameters) to a near equal size $49\mathrm{K}$ model ($124\mathrm{M}$ parameters) to isolate embedding capacity effects. We reduced the hidden dimension from $768$ to $648$ and the feed‑forward dimension from $2048$ to $1728$, while maintaining identical model depth which known to have a dominant impact on performance [tay2021, petty2023]. As shown in the table 3, the $49\mathrm{K}$ model incurs higher cross‑entropy loss and delivers inferior results on downstream benchmarks (ARC‑E, HellaSWAG, PIQA, and SCIQ) compared to its $24\mathrm{K}$ counterpart, indicating that simply expanding the embedding layer contributes far less to model expressivity than is often assumed.
@@ Appendix F Do the benefits of larger vocabularies persist with tied-embedding models?
Several studies show that tying input and output embeddings can change pre-training dynamics [Inan2017, lesci2025]. We therefore repeat our study with tied embeddings and report average frequent-word loss and cross-entropy on FineWeb-Edu in the table 4. The tied embedding variant shows broadly similar average frequent word loss and ce-loss to the untied baseline.
@@ Appendix H Learning rate exploration
We use a fixed learning rate of $6\times 10^{-4}$ for an $85\mathrm{M}$ non-embedding–parameter language model, mirroring the learning rate of GPT-3 with an equivalent non-embedding model size [gpt-3]. we further conducted a learning‑rate exploration with four additional learning rates: $2.4\times 10^{-3}$, $1.2\times 10^{-3}$, $1.5\times 10^{-4}$, $7.5\times 10^{-5}$. Table 6 and 7 report the resulting cross‑entropy loss and average word loss of frequent 2500 words in fineweb-edu across different vocabulary sizes and learning rates, using the same validation dataset described in Section 3.1.